OpenStudy (anonymous):

A particle moves on a surface whose equation is z = f(x, y) = (x − 1)^2+ y^2 Assume that the x and y coordinates of the particle at time t are given by x = 2 cos t and y = 2 sin t , where 0 ≤ t ≤ 2π. Use the total derivative rule to show that the rate of change of the height of the particle above the xy plane is given by dz/dt = 4 sin t. Hence find the maximum value of the height and the coordinates (x, y,z) of the particle when the maximum occurs.

OpenStudy (dumbcow):

\[\frac{dz}{dt} = \frac{dz}{dx}*\frac{dx}{dt} + \frac{dz}{dy}*\frac{dy}{dt}\] \[\frac{dz}{dx} = 2(x-1)\] \[\frac{dx}{dt} =-2\sin(t)\] \[\frac{dz}{dy} = 2y\] \[\frac{dy}{dt} = 2\cos(t)\] substitute into top equation to verify dz/dt = 4sin(t)

OpenStudy (anonymous):

Oh, it was simpler than I thought. Thanks!